6. DC Utilities, Inc., is trying to determine its cost of debt. The utility has a debt issue outstanding with 10 years to mat step by step solutions please. no excel solutions.
Answer #1

SOLUTION USING FINANCIAL CALCULATOR :

ANNUAL COUPON = 1000 X 6% = 60

N =10, PV = 1000 X 90% = 900, FV =1000

NOW HOW VALUES TO BE ENTERED IN CALCULATOR :

N =10, PV = 900, PMT = -60, FV = -1000

CPT I/Y = 7.45%

SO YTM = BEFORE TAX COST OF DEBT = 7.45%

AFTER TAX COST OF DEBT = YTM*(1-t) == 7.45%*(1-0.35) = 4.84%

Answer : (I) 7.45% (II) 4.84% (Thumbs up please)